A ray of light is incident on a surface of glass slab at an angle 450. If the lateral shift produced per unit thickness is 1/3 m, the angle of refraction produced is?

Grade:12th pass

1 Answers

Ravi
9 years ago
Lateral Shift(x)= t sin(i-r)/cos r
Given that lateral shift per unit thickness of slab= 1/3(= lateral shift/thickness=x/t)
x/l=1/3= Sin(45-r)/cos r
Solve the trig. eqns to get the value of r
….....................................................................................................................................